One of the educational sites on the Internet requires the creation of
User account protected with 8 characters password
Different Choose from letters: A, B, C, D, E, F
And numbers: 6 , 5 , 4 , 3 , 2 , 1, find the number of passwords that
They can be created in the following cases:
The password contains 4 numbers and 4 letters next to each other

Answer by ikleyn(52781)

.

(6*5*4*3) * (6*5*4*3) = 360 * 360 = 129600.         ANSWER

First multiplier  (6*5*4*3)  is for  4-letters block made of  6  different non-repeating letters.

Second multiplier  (6*5*4*3)  is for  4-digits block made of  6 different non-repeating digits.